# Sensitivity of $W$-boson measurements to low-mass right-handed neutrinos

### v1

@@ I Introduction
$+\frac{c_{NW}^{i}}{\Lambda^{2}}\bar{L}_{i}\sigma^{\mu\nu}\nu_{R}\sigma_{I}\tilde{W}_{\mu\nu}^{I}+{\textrm{h.c.}},$
where $c_{LNH}^{i}$, $c_{HNe}^{i}$, and $c_{NW}^{i}$ are effective couplings, $i=e,\mu,\tau$, the mass scale of the new particles is $\Lambda$, $L$ is the left-handed lepton doublet, $H$ is the Higgs doublet, and $\tilde{H}=i\sigma_{2}H^{*}$. Assuming real couplings, a non-zero $c_{HNe}^{i}$ causes a parity-violating V+A interaction, whereas the other two couplings cause chirality-flipping interactions.
@@
Here, we assume the associated Wilson coefficients are small enough for $\nu_{R}$ to dominantly decay outside the detector.
@@ II $W$ boson measurements
Measurements of the $W$ boson at the Tevatron and LHC [29, 30, 31, 32, 33, 34, 35, 36] are performed assuming that its production and decay follow the predictions of the Standard Model. Decays to right-handed neutrinos would affect these measurements, with a different impact at each collider due to differences between up-quark, down-quark, and sea-quark momentum distributions in the proton. We quantify the $c_{HNe}^{i}$ dependence of existing and potential measurements of the $W$ boson mass ($m_{W}$), cross-section, and cosine of the helicity angle in the $W$ boson decay.
@@ II.1 CDF $m_{W}$ measurement
The CDF measurement of the $W$ boson mass [34] was performed using template fits to distributions of final-state kinematic quantities projected transverse to the beamline. The templates were produced at different values of $m_{W}$, assuming SM production and decay. The measured mass is 76 MeV higher than the SM prediction, a difference that cannot be explained currently.
@@ II.2 ATLAS and CMS $m_{W}$ measurements
Figure 3: The production and decay of $W^{\pm}$ bosons at the LHC for the final state containing either $\nu_{L}$ or $\nu_{R}$. The $W^{\pm}$ boson momentum projected along the $z$-axis is in the direction of the proton, contributing a valence quark. The decay to $\nu_{L}$ is V-A, so the $\nu_{L}$ direction is opposite to that of the $W^{+}$ boson. In the case of $\nu_{R}$, the decay is V+A, and the direction is reversed.
@@ III Other constraints
A non-zero $c_{HN\ell}$ would impact various other phenomena ranging from lepton decays to supernova cooling. To compare limits arising from other measurements, we determine the effective mixing caused by $c_{HN\ell}$ between the active neutrinos and a fourth singlet neutrino. The leptonic $W$-boson couplings are
@@
Using this relation, Ref. [43] translated various mixing bounds to bounds on $c_{HNe}$. From their analysis, we conclude that various other experimental constraints are stronger than the bounds from $W$-boson measurements for neutrino masses higher than a few MeV.
For the $c_{HNe}$ coupling, the strongest bound transitions from $\pi\to\nu+e$ searches [44] interpreted as constraints on $U_{e4}$ [28] to nuclear $\beta$ decays [21]. In the notation of Ref. [21], $\tilde{\epsilon}_{L}=c_{HNe}v^{2}/2\Lambda^{2}$ and the $\beta$ decay constraints extend to masses up to typical nuclear reaction energies of a few MeV.
@@
Figure 7: 95% C.L. bounds on the Wilson coefficient for (a) electrons from $\beta$ and pion decay in orange and blue, respectively, and for (b) muons from muon and pion decay in orange and blue, respectively. The present constraint from the CMS measurement of $\Delta m_{W}=m_{W^{+}}-m_{W^{-}}$ is in green, while the projected LHC Run 3 constraints from lepton angular measurements are shown as the red dashed lines.
@@ IV Conclusions
Revisiting the CDF Run II analysis showed that the presence of a $W\to\ell\nu_{R}$ decay would shift the fitted $W$-boson mass upward by $76~\mathrm{MeV}$ for a Wilson coefficient near the perturbative bound, $c_{HN\ell}v^{2}/\Lambda^{2}\simeq 1.2$. Such a coefficient would also induce differences between measurements of $m_{W^{+}}$ and $m_{W^{-}}$ at ATLAS and CMS. No such deviations are observed, and a 95% CL bound of $|c_{HN\ell}|v^{2}/\Lambda^{2}<0.40$ is inferred, thereby excluding this interpretation of the CDF measurement.
@@ Acknowledgements.
CH and CP acknowledge the Senior IPPP Fellowship, which funded this work.

### v2

@@ I Introduction
$+\frac{c_{NW}^{i}}{\Lambda^{2}}\bar{L}_{i}\sigma^{\mu\nu}\nu_{R}\sigma_{I}\tilde{H}W_{\mu\nu}^{I}+{\textrm{h.c.}},$
where $c_{LNH}^{i}$, $c_{HNe}^{i}$, and $c_{NW}^{i}$ are effective couplings, $i=e,\mu,\tau$, the mass scale of the new particles is $\Lambda$, $L$ is the left-handed lepton doublet, $H$ is the Higgs doublet, and $\tilde{H}=i\sigma_{2}H^{*}$. Assuming real couplings, a non-zero $c_{HNe}^{i}$ causes a parity-violating V+A interaction, whereas the other two couplings cause chirality-flipping interactions  [^2] A renormalizable Yukawa term $Y^{i}\bar{L}_{i}\nu_{R}\tilde{H}$ (as well as the operator with coefficient $c_{LNH}^{i}$) generates neutrino mixing and produces the opposite-chirality interactions that we do not consider in our phenomenological study. This mixing is constrained to be small ($Yv/m_{N}\ll 1$), see e.g. [29, 54]. The $c_{NW}^{i}$ term is loop-generated in weakly coupled renormalizable UV completions, and is thus small in general..
@@ II $W$ boson measurements
Measurements of the $W$ boson at the Tevatron and LHC [30, 31, 32, 33, 34, 35, 36, 37] are performed assuming that its production and decay follow the predictions of the Standard Model. Decays to right-handed neutrinos would affect these measurements, with a different impact at each collider due to differences between up-quark, down-quark, and sea-quark momentum distributions in the proton. We quantify the $c_{HNe}^{i}$ dependence of existing and potential measurements of the $W$ boson mass ($m_{W}$), cross-section, and cosine of the helicity angle in the $W$ boson decay  [^3] The $W$ boson width is also affected by $c_{HNe}^{i}$, however its current value of $2014\pm 50$ MeV [59] corresponds to a 95% confidence-level constraint of $c_{HNe}^{i}<1.6$ for $\Lambda=v$ [15]. This is significantly weaker than the constraints we consider. Interpreting the recent ATLAS width measurement [32] in terms of $c_{HNe}^{i}$ would give $c_{HNe}^{i}=1.4\pm 0.6$, in agreement with the value we obtain from the CDF $m_{W}$ measurement but in tension with that obtained from LHC measurements of $m_{W^{+}}-m_{W^{-}}$..
@@ II.1 CDF $m_{W}$ measurement
The CDF measurement of the $W$ boson mass [35] was performed using template fits to distributions of final-state kinematic quantities projected transverse to the beamline. The templates were produced at different values of $m_{W}$, assuming SM production and decay. The measured mass is 76 MeV higher than the SM prediction, a difference that cannot be explained currently [39, 40, 41, 42, 43, 44, 45, 46].
@@ II.2 ATLAS and CMS $m_{W}$ measurements
Figure 3: The production and decay of $W^{\pm}$ bosons at the LHC for the final state containing either $\nu_{L}$ or $\nu_{R}$. The $W^{\pm}$ boson momentum projected along the $z$-axis is in the direction of the proton contributing a valence quark. The decay to $\nu_{L}$ is V-A, so the $\nu_{L}$ direction is opposite to that of the $W^{+}$ boson. In the case of $\nu_{R}$, the decay is V+A, and the direction is reversed.
@@ III Other constraints
A non-zero $c_{HN\ell}$ would impact various other phenomena, of which we explicitly consider pion decay, nuclear beta decay, and muon decay  [^4] We only discuss the most constraining measurements in the mass range 0.1–100 MeV. We have also investigated the constraint arising from $e^{+}e^{-}\to W^{+}W^{-}$ production at LEP [60], and find it to be weaker than the others ($|c_{HN\ell}|<0.51$).. To compare limits arising from these measurements, we determine the effective mixing caused by $c_{HN\ell}$ between the active neutrinos and a fourth singlet neutrino. The leptonic $W$-boson couplings are
@@
Using this relation, Ref. [54] translated various mixing bounds to bounds on $c^{i}_{HNe}$. From their analysis, we conclude that various other experimental constraints are stronger than the bounds from $W$-boson measurements for neutrino masses higher than a few MeV.
For the $c_{HNe}$ coupling, the strongest bound transitions from $\pi\to\nu+e$ searches [55] interpreted as constraints on $U_{e4}$ [29] to nuclear $\beta$ decays [22] near a neutrino mass of 0.2 MeV. In the notation of Ref. [22], $\tilde{\epsilon}_{L}=c_{HNe}v^{2}/2\Lambda^{2}$ and the $\beta$ decay constraints extend to masses up to typical nuclear reaction energies of a few MeV.
@@
[…]
@@ IV Conclusions
Revisiting the CDF Run II analysis showed that the presence of a $W\to\ell\nu_{R}$ decay would shift the fitted $W$-boson mass upward by $76~\mathrm{MeV}$ for a Wilson coefficient near the perturbative bound, $c_{HN\ell}v^{2}/\Lambda^{2}\simeq 1.2$. Such a coefficient would also induce differences between measurements of $m_{W^{+}}$ and $m_{W^{-}}$ at ATLAS and CMS. No such deviations are observed, and a 95% CL bound of $|c_{HN\ell}|v^{2}/\Lambda^{2}<0.27$ is inferred, thereby excluding this interpretation of the CDF measurement.
@@ Acknowledgements.
C.H. and C.P. acknowledge the Senior IPPP Fellowship, which funded this work. R.A. and M.S. are supported by STFC under Grants No. ST/X003167/1 and No. ST/X000745/1.